# Lifelong Graph Summarization with Neural Networks: 2012, 2022, and a Time Warp

### v1

@@
Abstract.
Summarizing web graphs is challenging due to the heterogeneity of the modeled information and its changes over time. We investigate the use of neural networks for lifelong graph summarization. Assuming we observe the web graph at a certain time, we train the networks to summarize graph vertices. We apply this trained network to summarize the vertices of the changed graph at the next point in time. Subsequently, we continue training and evaluating the network to perform lifelong graph summarization. We use the GNNs Graph-MLP and GraphSAINT, as well as an MLP baseline, to summarize the temporal graphs. We compare $1$-hop and $2$-hop summaries. We investigate the impact of reusing parameters from a previous snapshot by measuring the backward and forward transfer and the forgetting rate of the neural networks. Our extensive experiments on ten weekly snapshots of a web graph with over $100$M edges, sampled in 2012 and 2022, show that all networks predominantly use $1$-hop information to determine the summary, even when performing $2$-hop summarization. Due to the heterogeneity of web graphs, in some snapshots, the $2$-hop summary produces over ten times more vertex summaries than the $1$-hop summary. When using the network trained on the last snapshot from 2012 and applying it to the first snapshot of 2022, we observe a strong drop in accuracy. We attribute this drop over the ten-year time warp to the strongly increased heterogeneity of the web graph in 2022.
The source code and additional resources: https://github.com/jofranky/Lifelong-Graph-Summarization-with-Neural-Networks
@@ 1. Introduction
The idea of graph summarization is to generate a small representation $S$ of an input graph $G$, which preserves structural information necessary for a given task (Čebirić et al., 2019). Calculating the graph summary itself can be computationally expensive, especially for larger graphs (Blume et al., 2021). However, this reduction in size allows tasks, e.g., web data summary search (Konrath et al., 2012), entity retrieval (Ciglan et al., 2012), estimating cardinalities in database joins (Neumann and Moerkotte, 2011), and data visualization (Goasdoué et al., 2020), to be computed much faster than on the original graph (Blume et al., 2021). In machine learning, graph summaries are used to learn the embeddings for the original graph (Zhou et al., 2023) or for more efficient graph distillation (Gupta et al., 2023).
@@
The information preserved by a summary is defined by the summary model and considers the vertices’ $k$-hop neighborhoods. Thus, in a machine learning sense, graph summarization can also be expressed as a vertex classification task where each vertex in the graph belongs to a certain summary (Blasi et al., 2022), representing the class. This summarization operation has to be permutation invariant to the input, i. e., the vertices’ $k$-hop neighborhood (Blume et al., 2021; Konrath et al., 2012). Graph Neural Networks (GNNs) are by design permutation invariant to the input (Hamilton, 2020) and thus can predict the summary of each vertex (Blasi et al., 2022).
@@
Blasi et al. (Blasi et al., 2022) applied multiple GNNs to graph summarization. They considered graph summaries based on equivalence classes (Goasdoué et al., 2020) which are lossless with respect to specific features, such as the edge label of the outgoing edges of a vertex. The authors used the Dynamic Linked Data Observatory (DyLDO) dataset (Käfer et al., 2012) to compute the graph summary from the 6 May, 2012 snapshot. We extend the analysis of Blasi et al. by considering temporal graphs. We reuse the neural networks trained on snapshots from 2012 to summarize the snapshot from 2022. We call this a time warp.
@@
The problem with training a network only on the first snapshot is that the performance decreases when the graph changes over time (Zhou and Cao, 2021; Wang et al., 2020). This is especially true for graph summaries, where classes may appear or disappear between snapshots (Blasi et al., 2022). One simple solution is learning the summaries for each snapshot from scratch, resulting in a separate network per snapshot without any information about summaries of previous snapshots. However, this is not feasible due to the size of the snapshots in the millions of vertices. In this work, we address this problem by incrementally learning the network. We reuse the network trained on previous snapshots as a base for learning the summaries on the next snapshot. This training procedure is called lifelong graph learning (Chen and Liu, 2018).
@@
We first calculate the summaries of the vertices of each snapshot using a crisp algorithm (Blume et al., 2021). These summaries are used as the gold standard for training the GNNs. We continually train the networks on one snapshot after the other, i. e., we train one task after another. After training on a snapshot, we evaluate the network on the current, all past, and all future snapshots. We measure forward transfer, backward transfer, and forgetting rates of the networks based on classification accuracy (Lopez-Paz and Ranzato, 2017; Chaudhry et al., 2018). Forgetting is the performance drop of a network on snapshots trained on in earlier tasks after training on subsequent tasks (McCloskey and Cohen, 1989). Our results show that a network performs best on a snapshot $i$ when it is trained on the sequence of tasks up to that snapshot $i$. This observation is more prominent for the $2$-hop than the $1$-hop summary. After the ten-year time warp, reusing the network parameters from 2012 does not have a positive impact but also does not harm the performance in 2022. In summary, our contributions are:
@@
  • We extend graph summarization using GNNs from static graphs to temporal graphs. We experiment with two GNNs and a baseline multi-layer perceptron (MLP) on sequences of ten weekly snapshots from 2012 and 2022 of the DyLDO web graph.
@@
  • We show that a neural network performs best on a snapshot $i$ when being trained in a sequence of tasks up to and including that snapshot. The changes in the graph reduce the vertex classification performance. This observation is more prominent in 2022.
  • We show that an MLP is sufficient for a good classification performance in $2$-hop summaries, although it uses only $1$-hop information.
@@
  • Reusing parameters from a network trained on the last snapshot in 2012 has no benefit over a network trained from scratch on the first snapshot in 2022.
@@
The remainder of the paper is organized as follows: Below, we summarize related work. Section 3 introduces the neural networks and methods used for graph summarization. Section 4 describes the experimental apparatus. Our results are presented in Section 5. Section 6 discusses the results, before we conclude.
2. Related Work
First, we introduce the Graph Neural Networks (GNN) used in this work. Second, we look into recent approaches for graph summarization. Finally, we discuss literature on lifelong graph learning and pre-training.
2.1. Graph Neural Networks
A GNN uses the structural information of a graph contained in its edges and the features of the neighboring vertices to distinguish different vertices, e. g., to classify them. Many GNNs use a message-passing approach, e. g., the Graph Convolutional Network (GCN) (Kipf and Welling, 2017), the Graph Attention Network (Velickovic et al., 2018), and Graph Isomorphism Network (Xu et al., 2019).
GraphSAINT is a scalable GNN architecture using message-passing (Zeng et al., 2020). It is a sampling-based method that enables any GNN to be efficiently applied to large graphs. It uses a vertex-based, edge-based, or random-walk-based sampler to generate smaller subgraphs that are used to train the GNN.
Recently, Hu et al. (2021) introduced Graph-MLP, a GNN that does not use message passing. Graph-MLP consists of an MLP that is trained with a standard cross-entropy loss and a neighbor contrastive (NContrast) loss on the graph edges.
2.2. Graph Summarization
@@
FLUID (Blume et al., 2021) is a language and a generic algorithm for flexibly defining and efficiently computing graph summaries. It can express all existing lossless structural graph summaries w.r.t. the considered features such as the vertices’ edges, the vertex neighbors, and others. This computation of graph summaries is based on hash functions applied on a canonical order of the vertices features (Blume et al., 2021; Konrath et al., 2012). The application of neural networks as a hash function has mainly been considered in the context of security (Turčaník and Javurek, 2016; Lian et al., 2007) but recently also applied by Blasi et al. with different GNN architectures for summarization of statics graphs (Blasi et al., 2022). In their work, Blasi et al. used three GNN architectures based on message-passing (GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), and GraphSAINT (Zeng et al., 2020)), Graph-MLP (Hu et al., 2021) as a GNN without message-passing, and Bloom filter. A sampling method based on GraphSAINT is applied that considers the class distribution by sampling inverse to the number of occurrences of classes in the training set.
The work on static graphs of Blasi et al. is the foundation of our work. We extend it to consider temporal graphs in a lifelong learning setting, which we introduce below. We reuse the sampling strategy from Blasi et al. for training our networks and use GCN as the underlying GNN.
2.3. Lifelong Graph Learning
@@
Lifelong learning (Chen and Liu, 2018), also known as continual learning (Lesort et al., 2020), incremental learning (Chaudhry et al., 2018), and never-ending learning (Mitchell et al., 2018) is a learning procedure that imitates the lifelong learning ability of humans. It adds the importance of transferring and refining knowledge to the learned network. In lifelong learning (Fei et al., 2016), a network at time $t$ has performed a sequence of $t$ learning tasks $\mathcal{T}_{1},\mathcal{T}_{2},\dots,\mathcal{T}_{t}$ and has accumulated the knowledge $\mathcal{K}$ learned in these past tasks. At time $t+1$, it is faced with a new learning task $\mathcal{T}_{t+1}$. The network can use past knowledge to help performing the learning task $\mathcal{T}_{t+1}$.
@@
A lifelong learning model should be able to exploit previous knowledge to learn new tasks better or faster. Successful knowledge transfer occurs when the new tasks are similar to the previous ones (Ruvolo and Eaton, 2013). One issue of lifelong learning is catastrophic forgetting (McCloskey and Cohen, 1989). Catastrophic forgetting is the tendency of neural networks to abruptly and drastically forget previously learned information upon learning new information.
@@
Lifelong learning is also applicable to graphs. In lifelong graph learning, a challenge is that the graphs can grow or shrink over time. Adding and removing vertices in the graphs may also result in the inclusion of new class labels in classification tasks (Zhou and Cao, 2021; Wang et al., 2020). The main challenges of lifelong graph learning are discussed in recent surveys (Zhang et al., 2024; Febrinanto et al., 2022). These are the ability to detect new instances of a class or tasks in the graph, the ability to accommodate new knowledge in graphs based on additional data instances or classes, and the ability to perform graph representation learning efficiently and accommodate new knowledge by sampling the graph. We address the aspect of increasing the detection accuracy of already-seen classes and adding new classes to our networks.
An issue with vertex classification problems is that the class of a vertex depends on its edges and, thus, is dependent on its neighbors. Wang et al. (2022) convert vertex classification tasks into graph classification tasks, i. e., turn the problem from lifelong graph learning into a task on i.i.d. data.
Our approach of lifelong graph learning follows the approach of Galke et al. (2023). The authors continue training the GNN on a new snapshot of the graph once it is available. Among others, they compare reusing an existing network (warm restart) versus training a new network from scratch for each task (cold restart). The results show that the warm restart benefits the lifelong training performance.
3. Graphs and Summarization
@@ 3.1. Graphs
3.2. Classical Graph Summarization
@@
A graph summary maps a graph $G$ to a smaller representation $S$, which is a graph that preserves structural information of $G$ (Scherp et al., 2023). Each vertex of the original graph $G$ is a member of exactly one equivalence class (EQC) of the summary graph. An EQC represents a set of vertices that share common features, such as the same set of edge labels. We formulate graph summarization as a vertex classification task following Blasi et al. (2022). Our summary models depend on the $k$-hop neighborhood of a vertex. The $1$-hop summary model is the Attribute Collection ($M_{AC1}$) that Blasi et al. (Blasi et al., 2022) used. The $2$-hop summary model $M_{AC2}$ extends the $1$-hop model by also considering the outgoing edges of the neighbors of a vertex. The models only make use of outgoing edges and only differ in the size of the neighborhood that is considered for the summary. In $ACi$, we calculate the EQC of a vertex $x$ by calculating $h_{i}(x)$ that considers the $i$-hop neighborhood of the vertex. We set $h_{1}(x)=0$ and then, for $i\geq 1$,
$h_{i+1}(x)=\bigoplus_{(p,h_{i}(y))\in H_{i}(x)}\mathrm{hash}(p,h_{i}(y))\,,$
@@
where $H_{i}(x)=\{(p,h_{i}(y))\mid(x,p,y)\in E\}$ and $\bigoplus$ is the bitwise XOR operator. Note that $H_{i}(x)$ is a set and thus contains no duplicates. This is important since the XOR operation otherwise cancels out two values with identical hashes. The function $\mathrm{hash}(p,h_{i}(y))$ first concatenates $p$ and $h_{i}(y)$ to a string, which is then hashed. We use the default hashing function for strings in Python, which uses SipHash (Python, 2023; Aumasson and Bernstein, 2012).
@@
An illustration of summarizing a vertex is shown in Figure 1. Our model considers, at most, the $2$-hop neighborhood of a vertex. In the case of vertex $v_{1}$, we recursively consider edges in the $2$-hop neighborhood of vertex $v_{1}$. Hence, we consider each edge in Figure 1 except the edge $(v_{6},p_{2},v_{9})$.
Figure 1. $3$-hop neighborhood of a vertex $v_{1}$.
@@
We generalize this classical graph summarization by considering graphs that change over time. For a number of hops $i\in\{1,2\}$, the graph summary $S_{t}$ of a graph $G_{t}=(V_{t},E_{t},R_{t})$ at time $t$ is a tuple $(V^{S}_{t},E^{S}_{t},R^{S}_{t})$. $V^{S}_{t}=C\cup D$, where $C$ is the set of EQCs computed from the vertices $V_{t}$, and $D=\{d_{p,h_{i}(y)}\mid(x,p,y)\in E_{t}\text{ for some }x\}$ is a set of vertices further specifying the EQCs. These are the primary and secondary vertices of Blume et al. (Blume et al., 2021). $R^{S}_{t}$ is the set of predicates of the original graph $G_{t}$ that were considered for the summarization. $E^{S}_{t}$ are the outgoing edges of the summary vertices $V^{S}_{t}$, given by
$E^{S}_{t}=\{(v,(p,h_{i}(y)),d_{p,h_{i}(y)})\mid(x,p,y)\in E_{t}\text{ and }x\in C\}\,,$
where $d_{p,h_{i}(y)}\in D$.
3.3. Neural Graph Summarization
@@
When applying graph neural networks (GNNs) for graph summarization, we need to define the vertex features (Hamilton, 2020; Blasi et al., 2022). Instead of hashing, as above, the feature vector $h_{v}^{(0)}$ of a vertex $v$ is a multi-hot encoding of its outgoing edge labels, following Blasi et al. (Blasi et al., 2022). The motivation is that the multi-hot encoded feature vector can be easily extended with new elements once new predicates appear in the temporal graph. We note that multi-hot encoding is appropriate for graph summarization, as graph summary models only depend on the existence of edge or vertex labels, not their multiplicity (Blume et al., 2021).
@@
For the $1$-hop model $M_{AC1}$, the feature vector already contains all the information needed to calculate the summary, which is the label set of the vertex’s edges. For the $2$-hop model $M_{AC2}$, we consider the feature vector of $v_{1}$ and, in addition, consider the feature vectors of $v_{1}$’s neighbors. From these feature vectors, we know the labels of the outgoing edges of $v_{1}$ and its neighbors. Note that, in classical summarization, we also know which edge labels connect to which neighboring feature vectors. Here, we know the set of edge labels and the set of neighboring feature vectors but not which edge is associated with which neighbor. However, the neural networks can still differentiate between the equivalence classes in $M_{AC2}$.
@@
A solution to incorporate the edges is to represent an edge between a vertex and its neighbor as a vertex between them, labeled with the edge label multi-hot encoded as a feature vector. In this case, we use two hidden layers in the GNN to cover the additional hop. Where this procedure is not applied, we use one hidden layer. We evaluate both approaches, i. e., test whether a ($k-1$)-layer message passing can approximate a $k$-hop summary.
@@
For lifelong learning, we consider different snapshots of a graph $G$ associated with a timestamp $t$. At time $t$, we train a neural network to predict the summary $S_{t}$ of the graph $G$ at time $t$, which we refer to as task $\mathcal{T}_{t}$. We compute the number of EQCs present in the summary of time $t$ for each time $t$, and the number of added and deleted EQCs compared to time $t-1$. From this, we can also compute how many EQCs from time $t-1$ are still present at time $t$. We also count how many different EQCs in total have appeared in the snapshots $S_{1},\dots,S_{t}$, which corresponds to the size of the neural network’s output layer at time $t$.
3.4. Graph Neural Networks
@@
We use two GNNs in our study, following Blasi et al. (Blasi et al., 2022): GraphSAINT (Zeng et al., 2020), a classical message-passing model designed to scale with large graphs, and Graph-MLP (Hu et al., 2021), which is an alternative to the usual message-passing GNNs.
GraphSAINT (Zeng et al., 2020) uses a sampler (vertex-, edge-, or random walk-based) to create a sampled graph $G_{P}=(V_{P},E_{P},R_{P})$ from a graph $G$. A GCN model is dimensioned to fit the sampled graph. The embedding $h^{k}_{v}$ of each vertex $v$ in layer $k$ is calculated as
$h^{k}_{v}=\sigma\left(\sum_{u\in V_{P}}\tilde{A}_{P}[v,u](W^{k-1})^{T}h_{u}^{(k-1)}\right),$
where $\tilde{A}_{P}$ is the normalized and sampled adjacency matrix, $W$ is the weight matrix, and $\sigma$ is the activation function. We use GraphSAINT with jumping knowledge connections and either one or two hidden layers (Xu et al., 2018).. In contrast to the original networks, we follow Blasi et al. (2022) and do not use layer normalization because the summary of a vertex depends on the structural information of its neighborhood, i. e., each edge is equally important. Additionally, we ran our experiments measuring the influence of layer normalization (see Appendix C), which comes to the same conclusion.
Graph-MLP (Hu et al., 2021) consists of a two-layer MLP followed by a classifier layer. The first layer ${X^{1}}=\text{Dropout}({\sigma}(XW^{0}))$ uses a GELU-activation function $\sigma$ and dropout. In the second layer $Z=X^{1}W^{1}$, the NContrast loss $\mathrm{loss}_{\mathrm{NC}}$ is applied. The output of the last layer $Y=ZW^{2}$ is used for the vertex classification, where standard cross entropy loss $\mathrm{loss}_{\mathrm{CE}}$ is applied. The NContrast loss of the $i$-th vertex is formulated as:
$\ell_{i}=-\log\frac{\sum^{B}_{j=1}1_{[j\neq i]}\gamma_{ij}\exp(\mathrm{sim}(z_{i},z_{j})/\tau)}{\sum^{B}_{k=1}1_{[k\neq i]}\exp(\mathrm{sim}(z_{i},z_{k})/\tau)},$
where $\mathrm{sim}(z_{i},z_{k})$ denotes cosine similarity. $z_{i}$ is the vertex embedding of the vertex to be classified after the second layer $Z$, and $z_{k}$ is the embedding of a sample vertex after the second layer $Z$. $\tau$ denotes the temperature parameter and $\gamma_{ij}$ is the strength of the connection between vertices $i$ and $j$.
$\gamma_{ij}\begin{cases}\neq 0&\text{if $j$ is in the $n$-hop neighborhood of }i\\ =0&\text{if $j$ is not in the $n$-hop neighborhood of }i\,.\end{cases}$
The NContrast loss of the $i$-th vertex considers $r$-hop neighbors of the vertex as positive samples, while the other vertices are negative samples. In our experiments, we only consider the $1$-hop neighbors of a vertex as positive samples. The total NContrast loss is then $\mathrm{loss}_{\mathrm{NC}}=\frac{1}{B}\sum^{B}_{i=1}\ell_{i}\,$. The combined loss is calculated as $\mathrm{loss}_{\mathrm{final}}=\mathrm{loss}_{\mathrm{CE}}+\alpha\cdot\mathrm{loss}_{\mathrm{NC}}\,$. The weighting coefficient $\alpha$ is used to balance the two losses. For $\alpha=0$, it is an MLP using only cross entropy as loss function.
3.5. Sampling Procedure
@@
Due to the large size of our graphs, we use the sampler of Blasi et al. (2022) for training. The sampler is based on the GraphSAINT vertex sampler, with each vertex weighted according to the class distribution. This accounts for the imbalanced class distributions of the snapshots. The sampled subgraph of a vertex consists of either its $1$-hop neighborhood for $M_{AC1}$ or its $2$-hop neighborhood for $M_{AC2}$. The sampler is also applied to Graph-MLP to sample the $1$-hop neighborhoods. These are treated as positive samples. Vertices that are not connected via an edge are treated as negative samples.
One challenge of the sampling process when using the $2$-hop model is that the number of vertices can vary greatly between batches. We have to add all neighboring vertices of a sampled vertex since they determine the vertex’s class. However, the size of the neighborhood depends on the degrees of the sampled vertices. To avoid running out of memory, we restrict the sampler to a specific number of vertices in each batch as in Blasi et al. (2022).
4. Experimental Apparatus

### v2

# Lifelong Graph Learning for Graph Summarization

@@
Abstract
Summarizing web graphs is challenging due to the heterogeneity of the modeled information and its changes over time. We investigate the use of neural networks for lifelong graph summarization. After observing the web graph at a point in time, we train a network to summarize graph vertices. We apply this trained network to summarize the vertices of the changed graph at the next point in time. Subsequently, we continue training and evaluating the network to perform lifelong graph summarization. We use the GNNs Graph-MLP and GCN, as well as an MLP baseline, to summarize the temporal graphs. We compare $1$-hop and $2$-hop summaries. We investigate the impact of reusing parameters from a previous snapshot by measuring backward and forward transfer as well as forgetting rate. Our extensive experiments are on two series of ten weekly snapshots, from 2012 and 2022, of a web graph with over $100$M edges. They show that all networks predominantly use $1$-hop information to determine the summary, even when performing $2$-hop summarization. Due to the heterogeneity of web graphs, in some snapshots, the $2$-hop summary produces up to ten times as many vertex classes as the $1$-hop summary. When using the network trained on the last snapshot from 2012 and applying it to the first snapshot of 2022, we observe a strong drop in accuracy. We attribute this drop over the ten-year time warp to the strongly increased heterogeneity of the web graph in 2022. The source code and additional resources are available at https://github.com/jofranky/Lifelong-Graph-Summarization-with-Neural-Networks.
Index Terms:
@@ I Introduction
Graph summarization is the generation of a small representation $S$ of an input graph $G$, that preserves structural information necessary for a given task [1]. Calculating the graph summary itself can be computationally expensive [2]. However, the reduction in size allows tasks such as web data summary search [3] and data visualization [4], to be computed much faster than on the original graph [2]. In machine learning, graph summarization is used to improve the scalability of node embedding learning methods [5].
@@
The information preserved by a summary is defined by the summary model and considers the vertices’ $k$-hop neighborhoods. Thus, in a machine learning sense, graph summarization can be expressed as a vertex classification task where each vertex in the graph belongs to a certain summary [6], representing the class. This summarization operation has to be permutation invariant to the input, i. e., the vertices’ $k$-hop neighborhood [2, 3]. Graph Neural Networks (GNNs) are designed to be permutation invariant in this way [6].
@@
Blasi et al. [6] applied multiple GNNs to graph summarization. They considered graph summaries based on vertex equivalence classes [4] which are lossless with respect to specific features, such as the edge label of the outgoing edges of a vertex. The authors used the 6 May, 2012 snapshot of the Dynamic Linked Data Observatory (DyLDO) dataset [7]. We extend the analysis of Blasi et al. [6] with a temporal component by considering ten consecutive snapshots of DyLDO from each of 2012 and 2022. We evaluate reusing networks trained on 2012 for 2022, which we refer to as a time warp.
@@
The problem with training a network only on the first snapshot is that performance decreases when the graph changes over time [8]. This is especially true for graph summaries, where classes may appear or disappear between snapshots [6]. One simple solution is learning the summaries for each snapshot from scratch, resulting in a separate network per snapshot without any information about summaries of previous snapshots. In this work, we address this problem by incrementally learning the network. We reuse the network trained on previous snapshots as a base for learning the summaries on the next snapshot. This is called lifelong graph learning [9].
@@
We first calculate the summaries of the vertices of each snapshot using a crisp algorithm [2]. These summaries are used as the gold standard for training the GNNs. We continually train the networks on one snapshot after another. After training on a snapshot, we evaluate the network on that snapshot, and all past and future snapshots. We measure forward transfer, backward transfer, and forgetting rates of the networks based on classification accuracy [10, 11]. Forgetting is the performance drop of a network on snapshots trained on in earlier tasks after training on subsequent tasks [12]. Our results show that a network performs best on each snapshot when it is trained on the sequence of tasks up to and including that snapshot. This observation is more prominent for the $2$-hop than the $1$-hop summary. After the ten-year time warp, reusing the network parameters from 2012 neither improves nor harms the performance in 2022. In summary, our contributions are:
@@
  • We extend graph summarization using GNNs from static to temporal graphs. We experiment with two GNNs and a baseline MLP on sequences of ten weekly snapshots from 2012 and 2022 of the DyLDO web graph.
@@
  • We show that neural networks perform best on each snapshot when being trained on a sequence of tasks up to and including that snapshot. Changes in the graph reduce the performance which is more prominent in 2022.
  • We show that an MLP using only $1$-hop information is sufficient for $2$-hop summaries.
@@
  • Reusing parameters from a network trained in 2012 has no benefit over a network trained from scratch in 2022.
@@
The remainder of the paper is organized as follows: Below, we summarize related work. Section III introduces graph summarization. Section IV describes the experimental apparatus. Our results are presented in Section V. Section VI discusses the results, before we conclude.
II Related Work
II-A Graph Neural Networks
GNNs use the structural information of a graph contained in its edges and the features of the neighboring vertices to distinguish different vertices, e. g., to classify them. Many GNNs, including graph convolutional networks (GCNs) [13], graph attention networks (GANs) [14] and GraphSAINT [15] use message passing. GraphSAINT is a scalable GNN architecture that samples smaller subgraphs to enable any GNN to be trained on large graphs. Hu et al. [16] introduced Graph-MLP, a GNN that does not use message passing. Graph-MLP consists of a multi-layer perceptron (MLP) that is trained with a cross-entropy loss and a neighbor contrastive (NContrast) loss on the graph edges.
II-B Graph Summarization
@@
FLUID [2] is a language and a generic algorithm for flexibly defining and efficiently computing graph summaries. It can express all existing lossless structural graph summaries w.r.t. the considered features, such as the vertices’ edges. Computation of graph summaries is based on hash functions applied on a canonical order of the vertices features [2, 3]. The use of neural networks as a hash function has mainly been in the context of security [17, 18], but recently Blasi et al. [6] applied GCNs, GraphSAGE [19], GraphSAINT and Graph-MLP to summarizing static graphs, along with Blooom filters. We use a GraphSAINT-based sampling method by Blasi et al. [6], which considers the class distribution by sampling inverse to the number of occurrences of classes in the training set.
II-C Lifelong Graph Learning
@@
Lifelong learning [9] is a learning procedure that imitates the lifelong learning ability of humans. It adds the importance of transferring and refining knowledge to the learned network. In lifelong learning [20], a network at time $t$ has performed a sequence of $t$ learning tasks $\mathcal{T}_{1},\mathcal{T}_{2},\dots,\mathcal{T}_{t}$ and has accumulated knowledge from these past tasks. At time $t+1$, it is faced with a new learning task $\mathcal{T}_{t+1}$. The network can use past knowledge to help with learning task $\mathcal{T}_{t+1}$.
@@
A lifelong learning model should be able to exploit previous knowledge to learn new tasks better or faster. One issue in lifelong learning is catastrophic forgetting [12]. Catastrophic forgetting is the tendency of neural networks to drastically forget previous knowledge upon learning new information.
@@
In lifelong graph learning, a challenge is that the graphs can grow or shrink over time. Adding and removing vertices in the graphs may also result in the inclusion of new class labels in classification tasks [8]. We address the aspect of increasing the detection accuracy of already-seen classes and adding new classes to our networks. Galke et al. [21] compare reusing an existing network (warm restart) versus training a new network from scratch (cold restart). Their experiments show that warm restarts are generally preferred for lifelong graph learning.
III Graphs and Summarization
@@ III-A Labeled Graphs
III-B Classical Graph Summarization
@@
A graph summary maps a graph $G$ to a smaller representation $S$, which is a graph that preserves structural information of $G$ [24]. Each vertex of the original graph $G$ is a member of exactly one equivalence class (EQC) of the summary graph. Each EQC represents a set of vertices that share common features, such as the same set of edge labels. We formulate graph summarization as a vertex classification task following Blasi et al. [6]. Our summary models depend on the $k$-hop neighborhood of a vertex. We use the $1$- and $2$-hop versions of Attribute Collection, $M_{\mathrm{AC1}}$ and $M_{\mathrm{AC2}}$, respectively. $M_{\mathrm{AC1}}$ considers a vertex’s RDF properties and $M_{\mathrm{AC2}}$ also includes the neighbors’ properties. These models use of outgoing edges and only differ in the size of the neighborhood that is considered for the summary. In $M_{\mathrm{AC}k}$, we calculate the EQC of a vertex $x$ by calculating the following function $h_{k}(x)$ that considers the $k$-hop neighborhood of the vertex. We set $h_{1}(x)=0$ and then, for $k\geq 1$,
$h_{k+1}(x)\ \;=\!\!\bigoplus_{(x,p,y)\in E}\!\!\mathrm{hash}(p,h_{k}(y))\,,$
@@
where $\bigoplus$ is the bitwise XOR operator and the function $\mathrm{hash}(p,h_{i}(y))$ first concatenates $p$ and $h_{i}(y)$ to a string, which is then hashed. We use the default hashing function for strings in Python, which uses SipHash [25, 26].
@@
Figure 1 illustrates summarizing a vertex. Our model considers, at most, the $2$-hop neighborhood of a vertex. In the case of vertex $v_{1}$ (center), we recursively consider edges in the $2$-hop neighborhood of vertex $v_{1}$. Hence, we consider each edge in Figure 1 except the edge $(v_{6},p_{2},v_{9})$, which extends to the three-hop neighborhood.
Figure 1: $3$-hop neighborhood of a vertex $v_{1}$.
@@
We generalize this classical graph summarization by considering graphs that change over time. For a number of hops $i\in\{1,2\}$, the graph summary $S_{t}$ of a graph $G_{t}=(V_{t},E_{t},R_{t})$ at time $t$ is a tuple $(V^{S}_{t},E^{S}_{t},R^{S}_{t})$. $V^{S}_{t}=C\cup D$, where $C$ is the set of EQCs computed from the vertices $V_{t}$, and $D=\{d_{p,h_{i}(y)}\mid(x,p,y)\in E_{t}\text{ for some }x\}$ is a set of vertices further specifying the EQCs. These are the primary and secondary vertices of Blume et al. [2]. $R^{S}_{t}$ is the set of predicates of the original graph $G_{t}$. For $d_{p,h_{i}(y)}\in D$, the outgoing edges $E^{S}_{t}$ of the summary vertices $V^{S}_{t}$ are given by $E^{S}_{t}=\{(v,p,d_{p,h_{i}(y)})\mid(x,p,y)\in E_{t}\text{ and }x\in C\}$.
III-C Neural Graph Summarization
@@
When applying graph neural networks (GNNs) for graph summarization, we need to define the vertex features [27, 6]. Instead of hashing, as above, the feature vector $h_{v}^{(0)}$ of a vertex $v$ is a multi-hot encoding of its outgoing edge labels, following Blasi et al. [6]. The motivation is that the multi-hot encoded feature vector can be easily extended with new elements once new predicates appear in the temporal graph. We note that multi-hot encoding is appropriate for graph summarization, as graph summary models only depend on the existence of edge or vertex labels, not their multiplicity [2].
@@
For the $1$-hop model $M_{\mathrm{AC1}}$, the feature vector already contains all the information needed to calculate the summary, which is the label set of the vertex’s edges. For the $2$-hop model $M_{\mathrm{AC2}}$, we consider the feature vector of $v_{1}$ and, in addition, consider the feature vectors of $v_{1}\!$’s neighbors. From these feature vectors, we know the labels of the outgoing edges of $v_{1}$ and its neighbors. Note that, in classical summarization, we also know which edge labels connect to which neighboring feature vectors. Here, we know the set of edge labels and the set of neighboring feature vectors but not which edge is associated with which neighbor.
@@
An approach to incorporate the edges is to represent an edge between a vertex and its neighbor as a vertex between them, labeled with the edge label one-hot encoded as a feature vector. In this case, we use two hidden layers in the GNN to cover the additional hop. Where this procedure is not applied, we use one hidden layer. We evaluate both approaches, i. e., test whether message passing is necessary for a $2$-hop summary
@@
For lifelong learning, we consider different snapshots of a graph $G_{t}$ associated with a timestamp $t$. At time $t$, we train a neural network to predict the summary $S_{t}$ of the graph $G_{t}$ at time $t$, which we refer to as task $\mathcal{T}_{t}$. We compute the number of EQCs present in the summary of time $t$ for each time $t$, and the number of added and deleted EQCs compared to time $t-1$. From this, we can also compute how many EQCs from time $t-1$ are still present at time $t$. We also count how many different EQCs in total have appeared in the snapshots $S_{1},\dots,S_{t}$, which corresponds to the size of the neural network’s output layer at time $t$.
@@
We use two GNNs in our study: a GCN [13], a classical message-passing model, and Graph-MLP [16], which is an alternative to the usual message-passing GNNs.
@@
Due to the large size of our graphs, we use the sampler of Blasi et al. [6] for training. The sampler is based on the GraphSAINT vertex sampler, with each vertex weighted according to the class distribution. This takes into account the imbalanced class distributions of the snapshots.
III-D Complexity Analysis
Our approach has two steps: first, summarizing a graph w.r.t. a given summary model and, second, training the graph neural network. For the first step, we iteratively compute summaries of a snapshot-based temporal graph. This can be computed with an incremental, parallel algorithm in time $\mathcal{O}(N\,\Delta^{k})$ [28], where $N$ is the number of changes in the graph from one snapshot to the next, $\Delta$ is the maximum degree, and $k$ is the number of hops considered in the summary model – for us, $k\in\{1,2\}$. This applies to summary models definable in FLUID, excluding any optional preprocessing step involving inference. The approach is parallelized and scales to large graphs.
The complexity of the second step depends on the specific neural network, the sample of vertices $V^{\prime}\subseteq V$, number of layers $K$, and hidden dimension $d$. Training and inference time is $\mathcal{O}(Kd(|E|+|V|d))$ for GCN [13], training time is $\mathcal{O}(Kd(|E|+|V|d))$ and inference time is $\mathcal{O}(Kd)$ for Graph-MLP [16], and $\mathcal{O}(Kd)$ for MLP [16].
IV Experimental Apparatus